OpenStudy (anonymous):

Differential Equation: y'=10^(x+y)

OpenStudy (anonymous):

\[y=-\frac{ \ln(c-10^x) }{\ln10 }\] that's the answer

OpenStudy (anonymous):

@paki ?

OpenStudy (xapproachesinfinity):

hmm let's try

OpenStudy (xapproachesinfinity):

\[y'=10^{x+y}=10^x 10^y \Longrightarrow \frac{dy}{dx}=10^x10^y\] separation method leads to:\[\int \frac{dy}{10^y}=\int 10^xdx\]

OpenStudy (xapproachesinfinity):

try that!

OpenStudy (anonymous):

I did!

OpenStudy (xapproachesinfinity):

it is best if you write is : \[\int 10^{-y}dy=\int 10^xdx\]

OpenStudy (xapproachesinfinity):

hmm so what's the problem then?

OpenStudy (anonymous):

that I'm not smart enough to see what you see....

OpenStudy (anonymous):

not our professor nor the TA speak much the language tbh

OpenStudy (xapproachesinfinity):

well i'm not smart either, i don't see anything at this point

OpenStudy (anonymous):

so it's almost sign language

OpenStudy (anonymous):

from here it's 10^-y/ln10 right?

OpenStudy (xapproachesinfinity):

hmm i don't memorize the antiderivatives so my way would be writing that as natural exp which i know how to handle

OpenStudy (xapproachesinfinity):

the same goes for the right hand side

OpenStudy (anonymous):

you mean lne^10?

OpenStudy (xapproachesinfinity):

i meant \[10^{-y}=e^{\ln (10^{-y})}=e^{-y\ln10}\]

OpenStudy (xapproachesinfinity):

now i can do it easily with substitution

OpenStudy (xapproachesinfinity):

that's change of base

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

\[e^\ln(10^-x)=e^-\ln(10^x)\]

OpenStudy (xapproachesinfinity):

so we would have \[\int e^{-y\ln10}dy=\int e^{x\ln10}x\]

OpenStudy (xapproachesinfinity):

that suppoed to be dx not x

OpenStudy (anonymous):

times x?

OpenStudy (xapproachesinfinity):

no dx

OpenStudy (anonymous):

oh ok

OpenStudy (xapproachesinfinity):

gotta go keep trying :)

OpenStudy (xapproachesinfinity):

you should get that answer

OpenStudy (anonymous):

wait shouldn't I first take off the exponent?

OpenStudy (anonymous):

@xapproachesinfinity ?

OpenStudy (anonymous):

@freckles ?

OpenStudy (anonymous):

am I right?

OpenStudy (anonymous):

can I just take the exponent off?

OpenStudy (anonymous):

@zepdrix ?

OpenStudy (anonymous):

Please I'm trying this problem since yesterday

OpenStudy (freckles):

\[\int\limits_{}^{} a^x dx=\frac{a^x}{\ln(a)}+C ,a \in (0,1) \cup (1,\infty)\]

OpenStudy (freckles):

\[\int\limits_{}^{}10^{x}dx=\frac{10^x}{\ln(10)}+C\]

OpenStudy (freckles):

\[\int\limits_{}^{}10^{-y} dy \\ \text{ you could sub the } -y=u \\ -dy=du\]

OpenStudy (anonymous):

but here it's \[\int\limits e^(-yln10)\]

OpenStudy (anonymous):

sorry @freckles I'm not getting it

OpenStudy (anonymous):

@IrishBoy123 what am I missing here?

OpenStudy (phi):

can you integrate e^x dx ? you get back e^x what about integral e^(ax) dx where a is a constant ?

OpenStudy (anonymous):

1/a

OpenStudy (phi):

your problem is \[ \int e^{ax} \ dx\] except you are using y instead of x, and a= -ln(10)

OpenStudy (anonymous):

but it's ln, not e.......

OpenStudy (anonymous):

sorry I just don't follow the solutions, did you continue from where @xo_kansasprincess_xo finished?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@freckles I swear I don't get it

OpenStudy (anonymous):

and I ever thought about a problem for 2 days in a row

OpenStudy (anonymous):

never*

OpenStudy (phi):

I assume you can get to here \[ \int e^{-y\ln10}dy=\int e^{x\ln10} dx \] which we can write \[ \int e^{(-\ln10) y}dy=\int e^{(\ln10)\ x} dx \]

OpenStudy (phi):

ln(10) is just a number (a constant) and so is -ln(10)

OpenStudy (freckles):

\[ \text{ Let } y=a^x \\ \ln(y)=\ln(a^x) \\ \ln(y)=x \ln(a) \\ \frac{y’}{y}=\ln(a) \\ y’=y \ln(a) \\ \text{ recall } y=a^x \\ \text{ so we have} y’=a^x \ln(a) \\ \frac{dy}{dx}=a^x \ln(a) \\ dy=a^x \ln(a) dx \\ \text{ integrating both sides we have } \\ y+C=\int a^x \ln(a) dx \\ \text{ now } ln(a) \text{ is a constant multiple of that integral so we can bring it outside the integral } \\ y+C=\ln(a) \int a^x dx \\ \text{ now divide both sides by} \ln(a) \\ \frac{y+C}{\ln(a)}= \int a^x dx \\ \text{ recall again } y=a^x \\ \frac{a^x+C}{\ln(a)}=\int a^x dx \\ \frac{a^x}{\ln(a)}+\frac{C}{\ln(a)}=\int a^x dx \\ \frac{a^x}{\ln(a)}+K=\int a^x dx \\ \text{ where } K \text{ represents some contant } \]

OpenStudy (phi):

if you can integrate e^ax dx then you can do both sides of your problem

OpenStudy (freckles):

\[\int\limits 10^{-y}dy=\int\limits 10^xdx \] you can apply that formula I wrote on both sides of here or you can rewrite it and use a different method in which @phi is talking of

OpenStudy (freckles):

but either way you will still get the same result on both sides

OpenStudy (anonymous):

still I have e at the bottom

OpenStudy (freckles):

I don't see an e

OpenStudy (freckles):

unless you choose to write it that way

OpenStudy (anonymous):

if you do the integral

OpenStudy (phi):

after integrating , you can change the e^-ln10 y back to 10^-y

OpenStudy (freckles):

here is example \[\int\limits_{}^{}5^{-v} dv \\ \text{ \let } u=-v \\ du=-dv \\ -du=dv \\ \int\limits 5^{u} (-du)=- \int\limits 5^u du =-\frac{5^u}{\ln(5)}+C=-\frac{5^{-v}}{\ln(5)}+C\]

OpenStudy (anonymous):

I know that @freckles ..........

OpenStudy (freckles):

then what is the problem

OpenStudy (anonymous):

oh!

OpenStudy (freckles):

\[y'=10^{x+y} \\ \frac{dy}{dx}=10^x 10^y \\ \frac{dy}{10^{y}}=10^x dx \\ 10^{-y} dy=10^x dx \\ \text{ integrate both sides }\]

OpenStudy (anonymous):

and the 10? how do I expres the equation with y then?

OpenStudy (anonymous):

I get then \[-10^y=10^x+\ln10c\]

OpenStudy (phi):

just C, because a constant (ln10) times an arbitrary constant is still just an arbitrary constant.

OpenStudy (phi):

and the exponent should be -y (not +y)

OpenStudy (anonymous):

oh so 10^-y?

OpenStudy (phi):

-10^-y = 10^x + C

OpenStudy (freckles):

you can also do it that other way... \[\int\limits_{}^{}5^{-v} dv=\int\limits e^{\ln(5^{-v})} dv \text{ since} \ln(x) \text{ and } e^x \text{ are inverses } \\ =\int\limits e^{-v \ln(5)} dv \text{ by power rule for } \log \\ \text{ then do a sub } \\ \text{ \let } u=-\ln(5) v \\ du=-\ln(5) dv \\ \frac{-1}{\ln(5)} du =dv \\ \text{ so we have } \int\limits e^{-v \ln(5)} dv=\int\limits e^{u} \frac{-1}{\ln(5)} du=\frac{-1}{\ln(5)}e^u+C \\ =\frac{-1}{\ln(5)}e^{-\ln(5) v}+C\] \[=\frac{-1}{\ln(5)}e^{\ln(5^{-v})} +C \text{ by power rule again } \\ =\frac{-1}{\ln(5)}5^{-v}+C \text{ since again } \ln(x) \text{ and } e^x \text{ are inverse functions }\] just wanted to show you the other way in which was spoke here to integrate that one example

OpenStudy (anonymous):

oh ok

OpenStudy (freckles):

and i know you are past that point but I wanted you to see both ways so you can pick your favorite

OpenStudy (anonymous):

thanks! but I didn't understand how to keep on from here

OpenStudy (phi):

-10^-y = 10^x + C How about write it as 10^-y = C- 10^x now take the ln of both sides

OpenStudy (phi):

in other words, multiply both sides by -1 notice that -C can be written as C

OpenStudy (anonymous):

\[y=-\ln(lnC-10^x)\]

OpenStudy (anonymous):

I know C is constant though I try to show it as much as similar to the answer

OpenStudy (phi):

10^-y = C- 10^x ln (10^-y ) = ln ( C- 10^x )

OpenStudy (phi):

can you continue?

OpenStudy (anonymous):

I'm an idiot, that's ll I have to say

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

it's the middle of the night here, and other excuses

OpenStudy (anonymous):

thank you all so much!!!!!!!!!!!!!!!!!!!!!!!!!!11

OpenStudy (anonymous):

you were both very helpful, I'll give @freckles 'cause he's got less than you @phi but really thank you so much! I wish I could be helpful to you someday...

OpenStudy (anonymous):

good night fellas!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!111

OpenStudy (freckles):

Well I will become phi's fan since I already gave a medal to x approaches infinity guy.

OpenStudy (anonymous):

you helped me for such a long time, I don't take it for granted, you've brightened my day